OpenStudy (anonymous):

19.2 is what percent of 24? A. 12.5% B. 68% C. 76% D. 80%

myininaya (myininaya):

is means equals of means multiply what is your variable (I will call it n) percent = 1/100 Reading your question we get what? 19.2 is what percent of 24? 19.2 = n 1/100 * 24 19.2=n*1/100*24 19.2=24n/100 Solve this for n And then write n as a percentage.